OpenStudy (anonymous):

The prism has a surface area of 202 ft2. What would the surface area of the prism be if each dimension were tripled? ft2 Prism with a length of five feet, a width of nine feet, and a height of four feet

OpenStudy (anonymous):

@Lyrae @LoveMeLikeYouDo @Algorithmic @OregonDuck @EclipsedStar @sweetburger @LoveMeLikeYouDo

OpenStudy (anonymous):

plzz some one

OpenStudy (anonymous):

come on

OpenStudy (anonymous):

@welshfella @endlesslight254 @RosieF @timo86m @yomamabf @urbanmorgans @i0iz0gangster @onlinestudentxP @puppylover0617 @ashhhrodriguez @sweetburger @Durply @FutureVet @Godlovesme @HannahA @Jennjuniper @Kainui @Lyrae @Concentrationalizing @VampirePrincess @BlossomCake @nuccioreggie @mathstudent55

OpenStudy (anonymous):

halp plzz

OpenStudy (durply):

hi

OpenStudy (durply):

hello?

OpenStudy (durply):

@ItZzZDrAgOn

OpenStudy (durply):

Hello?

OpenStudy (anonymous):

yes

OpenStudy (durply):

1818

OpenStudy (durply):

medal plz

OpenStudy (vampireprincess):

is this 7th grade h.w /problem?

OpenStudy (anonymous):

how'd yu find it

OpenStudy (anonymous):

yes @VampirePrincess

OpenStudy (durply):

Old Surface Area = 202 ft^2 New Surface Area = 1818 ft^2 Notice we have this ratio: (new area)/(old area) = 1818/202 = 9 and how 3^2 = 9. This is no coincidence. The old surface area is multiplied by 3^2 = 9 to get the new surface area. If you asked "What would the surface area of the prism be if each dimension were quadrupled?", then you would multiply the old surface area by 4^2 = 16 to get the new surface area

OpenStudy (durply):

:)

OpenStudy (vampireprincess):

i agree with @Durply :)

OpenStudy (durply):

Medal plz :)

OpenStudy (anonymous):

how do i medal

OpenStudy (durply):

click best response on my

OpenStudy (durply):

answer

OpenStudy (durply):

Thx anything else?

OpenStudy (anonymous):

yea shure wait a minute…..

OpenStudy (durply):

i can do most just not graphs XD

OpenStudy (durply):

my weak point

OpenStudy (durply):

;)

OpenStudy (anonymous):

A prism has a surface area of 176 cm2. What is the surface area of a scaled image with a scale factor of 1 half?

OpenStudy (durply):

0.25

OpenStudy (anonymous):

how did u find it

OpenStudy (durply):

just find the square root of 1

OpenStudy (durply):

wait

OpenStudy (durply):

its 44

OpenStudy (durply):

176/4 = 44

OpenStudy (anonymous):

i thought so that my first answer

OpenStudy (anonymous):

hey one last one

OpenStudy (durply):

ok

OpenStudy (anonymous):

Which is true about the volume or surface area of these prisms? A. The surface area of B is greater than the surface area of A. B. The volume of A is greater than the volume of B. C. The volume of B is greater than the volume of A. D. The surface area of A is greater than the surface area of B.

OpenStudy (durply):

is there a picture

OpenStudy (anonymous):

yes

OpenStudy (durply):

im not sure but my answer would be A

OpenStudy (durply):

you there?

OpenStudy (anonymous):

yes

OpenStudy (durply):

is that all?

OpenStudy (anonymous):

do you have an xbox 360 and xbox live

OpenStudy (vampireprincess):

@ItZzZDrAgOn your in k12 also? and i would think the answer is A.

OpenStudy (durply):

um i have xbox i can get live

OpenStudy (durply):

ya im also k12

OpenStudy (durply):

what did you get?

OpenStudy (durply):

@ItZzZDrAgOn

OpenStudy (anonymous):

i got the same answer as you

OpenStudy (durply):

no what percent

OpenStudy (anonymous):

yea and if yu do get live my gamer tag is DRAGONBORN327

OpenStudy (anonymous):

yes i got 80%

OpenStudy (durply):

o ok nice

OpenStudy (durply):

i quickscope and play bo2 mostly

OpenStudy (anonymous):

yea me too and i gta v

OpenStudy (anonymous):

hence my bo2 pic

OpenStudy (durply):

ya lol

OpenStudy (durply):

ok ill go add you

OpenStudy (durply):

wait xbox 360?

OpenStudy (anonymous):

what yur gamer tag

OpenStudy (anonymous):

yes 360

OpenStudy (durply):

NmTs xSUPERIORx

OpenStudy (durply):

its long lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

do u have gtav

OpenStudy (durply):

nope

OpenStudy (anonymous):

ok

OpenStudy (durply):

see ya i play mc and satuff on pc most of the time but ill play bo2

OpenStudy (anonymous):

idk i just got bored of mine craft one day

OpenStudy (vampireprincess):

guys these days:)

OpenStudy (anonymous):

there are plenty of girls that play video games

OpenStudy (anonymous):

@VampirePrincess

OpenStudy (vampireprincess):

i know i am one of them too you know :)

OpenStudy (anonymous):

whats yur gamertag

OpenStudy (vampireprincess):

i do not have a gamertag i just play games . . . . alot sometimes

OpenStudy (vampireprincess):

like driving cars

OpenStudy (vampireprincess):

lame i am lame huh?